# neurips_data_2024.bib

### v1

@@ 5 Conclusion
In this work, we considered issues associated with current evaluation practices for dynamic link prediction in temporal graphs. To address computational limitations, edges in the test set are split into fixed-size batches, causing several issues: In continuous-time temporal graphs, fixed-size batches create varying-length time windows depend on the temporal activity patterns. For edges within a batch we further lose information on their temporal ordering. In discrete-time temporal graphs where snapshots are typically larger than the batch size, batches impose an ordering of edges which is not present in the data. Moreover, state-of-the-art approaches for dynamic link prediction have treated the batch size as a tunable parameter. However, changing the batch size actually changes the prediction task, resulting in incomparable results between different batch sizes.
We solve these issues by formulating the dynamic link forecasting task. Dynamic link forecasting acknowledges the resolution at which temporal interaction data is recorded and explicitly considers a forecasting horizon corresponding to a prediction time window of a fixed duration. Depending on the dataset and problem setting, the horizon may span seconds, minutes, hours, or longer, but crucially, time windows always span the same length. We evaluated dynamic link forecasting performance of nine state-of-the-art temporal graph learning approaches on 14 real-world datasets, comparing it to the common dynamic link prediction evaluation. We find substantial differences, especially for memory-based TGNNs. We provide data loader implementations for commonly used evaluation frameworks to facilitate practical applications of our evaluation approach.
@@
Limitations of our work include that our reformulation of the dynamic link prediction task suggests time-window-based approaches for model training, which however goes beyond the scope of our paper. After completing our experiments, we became aware of a current work proposing a correction technique that could account for some of the issues addressed by our work during training [su2024pres]. Future work could thus evaluate whether this correction approach can mitigate some of the differences observed in our results.
Despite these open issues, we believe that our work is an important and timely contribution that can help to improve evaluation practices in temporal graph learning.
Acknowledgments and Disclosure of Funding
@@ Appendix A Negative sampling approaches
[…]
Given a training set $E_{\text{train}}$ and test set $E_{\text{test}}$, each containing a sequence of edges $E_{t_{\text{min}}:t_{\text{max}}}$ in the temporal graph $G$, we can define the following commonly used sampling strategies for drawing negative samples $B_{i}^{-}$ for batch $B_{i}$ with $|B_{i}|=|B_{i}^{-}|$ [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu].
  • Random (rnd): Sample $B_{i}^{-}$ from $V^{2}$ without replacement. The subgraph corresponding to $B_{i}$ is assumed to be sparse, making it unlikely to sample a positive edge $e\in B_{i}$ as negative.
  • Historic (his): Sample $B_{i}^{-}$ without replacement from all training edges $E_{\text{hist}}=E_{\text{train}}\setminus\left\{(v_{j},u_{j})|t_{i\cdot b}\leq t_{j}\leq t_{(i+1)\cdot b}\right\}$ except the ones appearing at the same time as the edges in $B_{i}$. If $|E_{\text{hist}}|<|B_{i}|$, draw the remaining edges randomly as described above.
  • Inductive (ind): Sample $B_{i}^{-}$ without replacement from all unseen test edges $E_{\text{ind}}=E_{\text{test}}\setminus(E_{\text{train}}\cup\left\{(v_{j},u_{j})|t_{i\cdot b}\leq t_{j}\leq t_{(i+1)\cdot b}\right\})$ except the ones appearing at the same time as the edges in $B_{i}$. If $|E_{\text{ind}}|<|B_{i}|$, draw the remaining edges randomly as described above.
@@
Appendix B Datasets
@@
In this work, we use eight continuous-time and six discrete-time datasets, listed in Table 1. Here, we describe what systems were observed to create the datasets and plot the datasets’ link count histograms in Figure 5.
@@
(a) In the continuous-time temporal graphs, it is most common that at most one edge occurs per timestamp. Depending on the dataset, more than one edge per timestamp is more or less common.

### v2

@@ 5 Conclusion
In this work, we considered issues in current evaluation practices for dynamic link prediction in temporal graphs. To address computational limitations, edges are split into fixed-size batches, making the task “too easy” but at the same time “too hard”: “Too easy” because state-of-the-art approaches for dynamic link prediction have treated the batch size as a tunable parameter. Changing the batch size, however, changes the task difficulty, resulting in incomparable results between models using different batch sizes. In discrete-time temporal graphs, batches that include edges from the same timestamp cause information leakage that can be utilised by memory-based TGNNs. “Too hard” since in continuous-time temporal graphs, fixed-size batches create varying-length time windows leading to tasks of different difficulty for each duration, requiring the model to capture multiple temporal resolutions at once. Furthermore, for edges within a batch, information regarding their temporal ordering is lost. In discrete-time temporal graphs where snapshots are typically larger than the batch size, batches additionally impose an ordering of edges not present in the data.
We solve these issues by formulating the dynamic link forecasting task. Dynamic link forecasting acknowledges the resolution at which temporal interaction data is recorded and explicitly considers a forecasting horizon corresponding to a time window of fixed duration. Depending on the dataset and problem setting, the horizon may span minutes, hours, or longer, but crucially, time windows always span the same length. We evaluated the dynamic link forecasting performance of nine state-of-the-art temporal graph learning approaches on 14 real-world datasets, comparing it to dynamic link prediction. We found substantial differences, especially for memory-based TGNNs. We provide implementations for PyTorch Geometric and DyGLib to facilitate practical applications of our evaluation approach.
@@
Limitations of our work include that our reformulation of the dynamic link prediction task suggests time-window-based approaches for model training, which however goes beyond the scope of our paper. Furthermore, the metrics used in this work consider the problem of dynamic link prediction or forecasting as binary classification, using one negative sample for each positive edge. Other approaches (You et al., 2022) consider the problem as a ranking task (e.g. using the mean reciprocal rank (MRR)) and compare each positive edge against a large number of negative samples. These approaches typically use only one positive sample per batch which alleviates the information loss and provides a better estimate of the models’ precision due to the large number of negative samples. In contrast, our window-based approach allows the parallel processing of many positive samples in each time window, leading to a considerably faster evaluation by reasonably simplifying the task. Specifically, our simplifying assumption is that it is enough to make predictions at a lower resolution than available in the data, resulting in a trade-off between evaluation runtime and granularity of temporal information available to the TGNN models. For example, it may not be required to predict whether a customer will purchase a certain product within the next second; making such a prediction for the next day or week may be sufficient.
Impact Statement
There are no expected negative societal impacts that go beyond those of other foundational works in machine learning research.
Acknowledgements
@@ Appendix A Negative Sampling Approaches
Dynamic link prediction is typically framed as a binary classification problem to predict class 1 for existing links during a certain time window and 0 otherwise. Due to the sparsity of most real-world graphs, it usually suffices to train and evaluate using all existing (positive) edges and a sample of non-existing (negative) edges out of all possible edges $V^{2}$. In static link prediction, negative edges are typically sampled randomly from $V^{2}$ without replacement but Poursafaei et al. (2022) showed that this technique is ill-suited for dynamic link prediction. One reason is rooted in the characteristics of temporal graphs where already-seen interactions tend to repeat several times during the observation period. To address this issue, Poursafaei et al. (2022) introduced negative sampling which we cover in the following.
Given a training set $E_{\text{train}}$ and test set $E_{\text{test}}$, each containing a sequence of edges $E_{t_{\text{min}}:t_{\text{max}}}$ in the temporal graph $G$, we can define the following commonly used sampling strategies for drawing negative samples $B_{i}^{-}$ for batch $B^{+}_{i}$ with $|B^{+}_{i}|=|B_{i}^{-}|$ (Poursafaei et al., 2022; Yu et al., 2023).
  • Random: Sample $B_{i}^{-}$ from $V^{2}$ without replacement. The subgraph corresponding to $B^{+}_{i}$ is assumed to be sparse, making it unlikely to sample a positive edge $e\in B_{i}$ as negative.
  • Historic: Sample $B_{i}^{-}$ without replacement from all training edges $E_{\text{hist}}=E_{\text{train}}\setminus\left\{(v_{j},u_{j})|t_{b\cdot i}\leq t_{j}\leq t_{b\cdot(i+1)}\right\}$ except the ones appearing at the same time as the edges in $B^{+}_{i}$. If $|E_{\text{hist}}|<|B^{+}_{i}|$, draw the remaining edges randomly as described above.
  • Inductive: Sample $B_{i}^{-}$ without replacement from all unseen test edges $E_{\text{ind}}=E_{\text{test}}\setminus(E_{\text{train}}\cup\left\{(v_{j},u_{j})|t_{b\cdot i}\leq t_{j}\leq t_{(i+1)\cdot b}\right\})$ except the ones appearing at the same time as the edges in $B^{+}_{i}$. If $|E_{\text{ind}}|<|B^{+}_{i}|$, draw the remaining edges randomly as described above.
@@
Appendix B Normalized Mutual Information
Normalized mutual information is an information-theoretic measure based on mutual information. It is based on mutual information, which for two random variables $X$ and $Y$ captures the bits of information we gain about the outcome of $Y$ if we know the outcome of $X$ and vice-versa. A formal definition is given in the following:
Definition B.1 (Mutual Information).
Consider two random variables, $X$ and $Y$ with joint probability mass function $p\left(x,y\right)$ and marginal probability mass functions $p\left(x\right)$ and $p\left(y\right)$. Mutual information is the reduction in the uncertainty of $X$ due to the knowledge of $Y$ defined as (Cover & Thomas, 2006)
$I\left(X,Y\right)=\sum_{x\in X}\sum_{y\in Y}p\left(x,y\right)\log\frac{p\left(x,y\right)}{p\left(x\right)p\left(y\right)}.$
Note that mutual information can be defined using different logarithms. The intuitive understanding described above using bits of information is defined using $\log_{2}$. This work utilizes the implementation of the Python library scikit-learn (Pedregosa et al., 2011) which uses the natural logarithm $\log_{e}$.
The specific value of mutual information depends on the entropy
$H(X)=-\sum_{x\in X}p(x)\log p(x)$
of the underlying random variables, and is thus difficult to compare across different settings. To address this issue, normalized mutual information provides a measure between zero and one that is normalized based on the entropies of the underlying random variables (Vinh et al., 2010). We use the following normalization as implemented in scikit-learn’s function normalized_mutual_info_score:
$\text{NMI}(X,Y)=\frac{I(X,Y)}{\frac{1}{2}\cdot(H(X)+H(Y))}$
In the context of our work, we use the NMI to capture the loss of temporal information. In Figure 2, we use the NMI to measure how much information about the edges’ timestamps is lost by grouping the edges into batches. Specifically, we use the number $i$ of each batch $B_{i}^{+}\cup B_{i}^{-}$ of the definition of dynamic link prediction in Section 2 assigned to each edge in the corresponding batch as one random variable and the timestamps of the edges as the other. With this setup, we can measure how much information about the timestamps of edges we gain – or keep – based on the batch number only.
Additionally, we use the NMI in Table 2 to quantify the difference between the assignments of edges to time windows and batches respectively. Similar as above, we assign to each edge in the test set of each dataset its batch number and also a time window number corresponding to the time window the edge belongs to and then compute the NMI between those two variables.
In the following, we present an educational example that illustrates the usage of NMI in the context of our work to facilitate comprehension among readers who may not be familiar with the measure:
Example B.1.
Let $G$ be a temporal graph with nodes $V=\left\{a,b,c\right\}$ and temporal edges
$E=\left((a,b,1),(b,c,2),(c,a,2),(a,c,4),(a,b,5),(b,a,5)\right).$
We define the timestamps $t_{i}$ of the edges $e_{i}\in E$, the batch number for $b=2$ and the window number with $h=1$ as random variables $X,Y$ and $Z$, respectively. We obtain the following realizations of $X,Y$ and $Z$:
$\mathbf{x}$ $=(1,2,2,4,5,5)$
$\mathbf{y}$ $=(1,1,2,2,3,3)$
$\mathbf{z}$ $=(1,2,2,3,4,4)$
In the next step, we compute the marginal probability mass functions $p(X=x),p(Y=y)$ and $p(Z=z)$:
$p(X=1)$ $=\frac{1}{6},$ $p(X=2)$ $=\frac{2}{6},$ $p(X=4)$ $=\frac{1}{6},$ $p(X=5)$ $=\frac{2}{6},$
$p(Y=1)$ $=\frac{2}{6},$ $p(Y=2)$ $=\frac{2}{6},$ $p(Y=3)$ $=\frac{2}{6},$
$p(Z=1)$ $=\frac{1}{6},$ $p(Z=2)$ $=\frac{2}{6},$ $p(Z=3)$ $=\frac{1}{6},$ $p(Z=4)$ $=\frac{2}{6}.$
Afterward, we compute the joint probability mass functions $p(X=x,Y=y)$, $p(X=x,Z=z)$, and $p(Y=y,Z=z)$:
$p(X=1,Y=1)$ $=\frac{1}{6},$ $p(X=2,Y=1)$ $=\frac{1}{6},$$\dots$
$p(X=1,Z=1)$ $=\frac{1}{6},$ $p(X=2,Z=2)$ $=\frac{2}{6},$$\dots$
$p(Y=1,Z=1)$ $=\frac{1}{6},$ $p(Y=1,Z=2)$ $=\frac{1}{6},$$\dots$
Using these marginal and joint probabilities, we compute the mutual information $I(X,Y)$, $I(X,Z)$, and $I(Y,Z)$ using the formula from above:
$I(X,Y)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{1}{6}\cdot\frac{2}{6}}+\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 0.868$
$I(X,Z)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{1}{6}\cdot\frac{1}{6}}+\frac{2}{6}\log_{e}\frac{\frac{2}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 1.330$
$I(Y,Z)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{1}{6}}+\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 0.868$
Finally, we obtain the normalized mutual information (NMI) by normalizing based on the entropies of both random variables:
$\text{NMI}(X,Y)$ $=\frac{0.868}{\frac{1}{2}(1.330+1.099)}\approx 0.715$
$\text{NMI}(X,Z)$ $=\frac{1.330}{\frac{1}{2}(1.330+1.330)}\approx 1.0$
$\text{NMI}(Y,Z)$ $=\frac{0.868}{\frac{1}{2}(1.099+1.330)}\approx 0.715$
Note that the NMI calculations in the main part of the paper are done accordingly. Specifically, the NMI values of the bottom-most continuous-time dataset and the right-most discrete-time dataset in Figure 1 are calculated as demonstrated with $\text{NMI}(X,Z)$, while the other NMI values in the figure as well as all values in Figure 2 are calculated following the example of $\text{NMI}(X,Y)$. Table 2 reports NMI values that follow the example calculation of $\text{NMI}(Y,Z)$.
Appendix C Datasets
@@
In this work, we use eight continuous-time and six discrete-time datasets, listed in Table 1. In the following, we describe what systems were observed to create the datasets and analyse the data in more detail.
C.1 Descriptions
We list details about each dataset in the following:
  • Enron (Shetty & Adibi, 2004) is a bipartite continuous-time graph where nodes are users and the temporal edges represent emails sent between users. Emails with multiple recipients are recorded as separate and simultaneously occurring edges, one per recipient. The temporal edges are resolved at the second level and the dataset spans approximately 3.6 years.
  • UCI (Panzarasa et al., 2009) is a unipartite continuous-time social network dataset from an online platform at the University of California at Irvine. The nodes represent students and the timestamped edges represent communication between the students. The dataset spans approximately six and a half months.
  • MOOC (massive open online course) (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent users and units in an online course, such as problems or videos. Temporal edges are resolved at the second level and encode when a user interacts with a unit of the online course. The dataset spans approximately one month.
  • Wikipedia (Wiki.) (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent editors and Wikipedia articles. The timestamped edges are resolved at the second level and represent when an editor has edited an article. The dataset spans approximately one month.
  • LastFM is a bipartite continuous-time graph where nodes represent users and songs. Temporal edges are resolved at the second level and model the users’ listening behavior and represent when a user has listened to a song. The dataset was originally published by Celma (2010) and later filtered by Kumar et al. (2019) for use in a temporal graph learning context.
  • Myket (Loghmani & Fazli, 2023) is a bipartite continuous-time graph where nodes represent users and Android applications. The timestamped edges represent when a user installed an application. The dataset spans approximately six and a half months.
  • Social Evolution (Social) (Madan et al., 2012) is a unipartite continuous-time graph of the proximity between the students in a dormitory, collected between October 2008 and May 2009 using mobile phones. Temporal edges connect students when they are in proximity and are resolved in seconds.
  • Reddit (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent Reddit users and their posts. The timestamped edges are resolved in seconds and represent when a user has made a post on Reddit. The dataset spans approximately one month.
  • UN Vote (UN V.)(Voeten et al., 2009; Poursafaei et al., 2022) is a weighted unipartite discrete-time graph of votes in the United Nations General Assembly between 1946 and 2020. Nodes represent countries and edges connect countries if they both vote “yes”. The dataset is resolved at the year level and edge weights represent how many times the two connected countries have both voted “yes” in the same vote.
  • US Legislators (US L.) (Huang et al., 2020; Fowler, 2006; Poursafaei et al., 2022) is a weighted unipartite discrete-time graph of interactions between legislators in the US Senate. Nodes represent legislators and edges represent co-sponsorship, i.e., edges connect legislators who co-sponsor the same bill. The dataset is resolved at the congress level and edge weights encode the number of co-sponsorships during a congress.
  • UN Trade (UN Tr.) (MacDonald et al., 2015; Poursafaei et al., 2022) is a directed and weighted unipartite discrete-time graph of food and agricultural trade between countries where nodes represent countries. The dataset spans 30 years and is resolved at the year level. Weighted edges encode the sum of normalized agriculture imports or exports between two countries during a given year.
  • Canadian Parliament (Can. P.) (Huang et al., 2020; Poursafaei et al., 2022) is a weighted unipartite discrete-time political network where nodes represent Members of the Canadian Parliament (MPs) and an edge between two MPs means that they have both voted “yes” on a bill. The dataset is resolved at the year level and the edges’ weights represent how often the two connected MPs voted “yes” on the same bill during a year.
  • Flights (Schäfer et al., 2014; Poursafaei et al., 2022) is a directed and weighted unipartite discrete-time graph where nodes represent airports and edges represent flights during the COVID-19 pandemic. The edges are resolved at the day level and their weights are given by the number of flights between two airports during the respective day.
  • Contacts (Cont.) (Sapiezynski et al., 2019; Poursafaei et al., 2022) is a unipartite discrete-time proximity network between university students. Nodes represent students who are connected by an edge if they were in close proximity during a time window. The dataset is resolved at the 5-minute level and spans one month.
C.2 Edge Activity patterns
Similar to Figure 3, we show the link densities over time for all datasets in Figure 5. While some datasets such as Enron, UCI, and LastFM contain very inhomogeneously distributed activity, others exhibit reoccurring or seasonal patterns, e.g. MOOC, Wikipedia, and Reddit. In addition to seasonality, we also observe trends like a growing number of edges in the Myket, UN Vote, and UN Trade datasets, which are also often found in standard time series data (Brockwell & Davis, 2016).
(a) Continuous-time temporal graphs resolved in seconds. For visualisation purposes, edges are binned into 6-hour time periods and then represented as one bar in the histogram.
(b) Discrete-time temporal graphs where each bar in the histogram corresponds to one snapshot. Since the snapshots capture different time ranges for each dataset (see Table 1), each dataset has its own temporal resolution.
Figure 5: Real-world datasets exhibit diverse edge occurrence patterns that are visualised using the edge density across time, i.e., histograms counting the number of edges per timestamp. Dashed lines divide the datasets into 70% train, 15% validation, and 15% test sets as used in Section 4.
C.3 Time Window Durations for Different Batch Sizes
In Figure 6 we evaluate the dependency between batch size and time window for empirical temporal graphs. The figure shows the distribution of batch durations (y-axis) for different batch sizes (x-axis). We observe that, both in continuous- and discrete-time temporal graphs, a single batch size can create time windows with varying durations even within the same dataset. For continuous-time temporal graphs, we typically have much bigger batches than edges per timestamp such that the time window defined by the batches becomes long (cf. Table 1). The number of edges per snapshot in discrete-time temporal graphs is generally larger than the batch size $b$ in any period regardless of the density (Table 1). This means that edges in a batch often belong to the same snapshot leading to small window durations.
Figure 6: Using a low opacity value for individual points, the distribution of time window durations is shown for different batch sizes. I.e. points appear less see-through with an increasing number of points with the same duration and batch size stacked on top of each other.
C.4 Number of Links per Timestamp
In Figure 7, we show how often each number of links with the same timestamp (x-axis) occurs in each dataset. We can observe that all continuous-time datasets have a large number of edges that appear at a unique point in time. Additionally, the number of edges that occur at once for many continuous-time datasets is less than ten (MOOC, Wiki., Myket, Reddit) and less than 50 for most datasets except for Enron which has outliers with more than 1000 links at a certain timestamp. For discrete-time datasets, many edges occur in each snapshot.
@@ C.4 Number of Links per Timestamp
(a) In the continuous-time temporal graphs, it is most common that at most one edge occurs per timestamp.